The weight of the potatoes is approximately normally distributed with population mean μ=10 ounces and population standard deviation σ=1.5 ounces. Use 68-95-99.7 rule to answer the questions below:

a). What is the probability that a randomly selected potato weighs over 13 ounces?

b). What is the probability that a randomly selected potato weighs below 8.5 ounces?

c). What is the probability that a randomly selected potato weighs between 8.5 ounces and 10 ounces?

d).What is the probability that a randomly selected potato weighs between 8.5 ounces and 13 ounces?
Answer #1


Solution :

Given that,

Using Empirical rule,8.5 < X <

P(\mu - 1\sigma< X < \mu + 1\sigma) = 68%

P(\mu - 2\sigma< X < \mu + 2\sigma) = 95%

P(\mu - 3\sigma< X < \mu + 3\sigma) = 99.7%

(a)

P(X > 13) = 1 - 0.975 = 0.025

(b)

P(X < 8.5) = 0.16

(c)

P(8.5 < X < 10) = P(X < 10) - P(X < 8.5) = 0.5 - 0.16 = 0.34

(d)

P(8.5 < X < 13) = P(X < 13) - P(X < 8.5) = 0.975 - 0.16 = 0.815